One gram mole of a monoatomic perfect gas λ =5/3 at 27° C is adiabatically compressed from an initial pressure of 1 atm to final pressure of 50 atm. Calculate 

the resulting final temperature.


1
Expert's answer
2021-04-13T11:10:28-0400

To be given in question

"T_{1}=27\u00b0=300K"

"Pressure P_{1}=1atm"

"Pressure P_{2} =50atm"

"\gamma=\frac{5}{3}"

To be asked in question

"T_{2}=?"

We know that

"PT^\frac{\gamma}{\gamma-1}. =constant"

"(\frac{P_{1}}{P_{2}})^{1-\gamma}=(\frac{T_{2}}{T_{1}})^{\gamma}"

Put value

"(\frac{1}{50})^{1-\frac{5}{3}}=(\frac {T_{2}}{300})^\frac{5}{3}"

"(\frac{1}{50})^{-\frac{2}{3}}=(\frac {T_{2}}{300})^\frac{5}{3}"

"({50})^{\frac{2}{3}}=(\frac {T_{2}}{300})^\frac{5}{3}"


"T_{2}=1434.52K"

"T_{2}=1161\u00b0C"